To: J. Cook, D. Helms, W. Skinner
cc: Division (M7), FEL Coordination Group
From: F. Dylla
Subject: IR Demo Project Weekly Report, February 23-27, 1998
Date: February 27, 1998
Highlights for the week include: delivery of the first 180 bend
vacuum chamber to the FEL vault ready for installation within
the east arc bend magnet; successful installation of key components
of the downstream optical cavity; refurbishment of the cathode
ball and manufacture of getter shields that are needed to re-assemble
the FEL gun; held the first class for training FEL operators.
Discussions were held on Feb. 23 with Jim Murday, head of the
Chemistry Division at NRL about the Division's interest in sponsoring
surface chemistry experiments with the FEL.
Eric Rohlfing from DOE-BES has forwarded a request to us for comparative
FEL specifications (vs. conventional lasers) that he will use
for discussions with the chemical physics and combustion science
community who are meeting at the ACS meeting in Dallas in late-March.
The SF6 recovery system was received and installed this week.
This will reduce our operating costs by about $7500 for SF6 every
time the gun cathode is replaced or the gun is otherwise refurbished.
The first 180° bend vacuum chamber was instrumented and cleaned.
It is in the FEL vault awaiting installation in the 3F region.
The installation of a Laser Safety System (LSS) for User Lab Room
4 is in progress to support the production of polarized He3 targets
for Hall A. This is the second FEL User Lab that has been made
available for Nuclear Physics activities involving Class 4 lasers.
In addition to providing essential services to the Nuclear Physics
program, the activity has enabled us to implement a smart card
access control system for the User Labs.
The downstream pellicle and the LSS shutter were installed in
the optical cavity and tested this week.
Testing of cavity microphonics, RF gradient and phase stability
was done this week. A cavity tuning algorithm (P-Tune) is scheduled
for this evening.
Miram curves for the 8 kW cryomodule klystrons were taken this
week at cathode voltages of 11.6 and 14.0 kV. Their heater voltages
were adjusted to extend their operating lifetimes at 14.0 kV.
An engine hoist was modified to fit under the optical tables and
certified to lift 1800 lb. It will be used to install or remove
the wiggler or other components of similar weight which can not
be reached with a fork lift.
The valve stem and bellows on the cryounit's J-T valve was replaced
this week to fix a small helium leak.
Gun refurbishment is proceeding, though a bit slower than we had
at first envisioned. Polishing of the ball is complete. The
stainless steel shield for the getter pumps is in fabrication
at a local contractor and will be delivered either late today
(27 Feb. 98) or early tomorrow. It will likely be cleaned over
the weekend and installed early next week, after which the gun
will be buttoned up and baked. A detailed plan has been worked
out for readying the gun for operation. We project gun operations
to resume on swing shift Sunday, 15 Mar 98.
An immediate impact of the gun delay is the need to restructure
our commissioning plans after turn-on to expedite GEN target irradiation.
Consequently, we will run tune-up beam to the straight-ahead
dump as quickly as possible, then insert the target chamber.
Presumably installation of the GEN target will occur on or about
23 Mar 98, with deinstallation at the end of that week. We will
defer all 60 pC setup until after the GEN targets are prepared.
There has been a general slippage in various installation tasks
due to higher-priority needs on CEBAF. A plus in the gun delay
is that the optical cavity may be completed by lock-up (provided
modest Laboratory support can be brought to bear on related installation
tasks), as will the MPS and probably installation of the final
air-core correctors. Therefore, following GEN, we should be able
to go after first light unabated (unforeseen technical difficulties
A "checkout" of the two multislit transverse-emittance diagnostics was done via simulation (PARMELA) with the latest injector settings for 60 pC. The basic findings are the following:
- At the lower emittances associated with 60 pC (vs. 135 pC on which the design of the multislits was baselined), the multislits still perform well.
- Multislit #1 (located between the first and second quads of the telescope following the cryounit) is sensitive to (1) RF phase of the buncher cavity, (2) magnetic field of gun solenoid #1 (emittance) and solenoid #2 (alpha), and (3) RF phase of the cryounit's entrance cavity.
- Multislit #2 (located at the injection point just prior to the cryomodule) is sensitive to the RF phase of the cryounit's exit cavity.
These findings verify that it is feasible to define a procedure
for using the multislits to identify and isolate drifts from nominal
settings in key components of the injector.